OpenStudy (anonymous):

how do you solve a parabola? y=x^2-4x+3 i know you find the intercepts, y=3 and x=3, 1. but what next?

OpenStudy (anonymous):

what do you want done? what do you mean to "solve a parabola"?

OpenStudy (anonymous):

like, graph a parabola... i think

OpenStudy (anonymous):

you want to find the axis of symmetry and the vertex

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

in the form \[ax^2+bx+c\] the axis of symmetry is \[x=\frac{-b}{2a}\] the vertex is \[(\frac{-b}{2a}, f(\frac{-b}{2a}))\]

OpenStudy (anonymous):

so y=x^2-4x+3 would become x=4/2? for the axis of symmetry

OpenStudy (anonymous):

and for the vertex it would be 4/2 times f(4/2)? what does the f mean?

OpenStudy (anonymous):

as in f(x) plug in -b/(2a) into the equation of the parabola to get the y-coordinate

OpenStudy (anonymous):

so vertex is at (2, -1)

OpenStudy (anonymous):

it is a point

OpenStudy (anonymous):

put 4/2 (-b/2a) where f(x) is?

OpenStudy (anonymous):

no.. f(x)=x^2-4x+3 this is the equation of your parabola the axis of symmetry is the vertical line that passes through the x-coordinate of your vertex To get the y-coordinate, plug in the x-coordinate into f(x) so f(2)=-1

OpenStudy (anonymous):

where'd you get f(2)=-1?

OpenStudy (anonymous):

plug 2 into f(x) f(2)=2^2-4(2)+3=4-8+3=-1

OpenStudy (anonymous):

oh ok alright so the f(4/2) = 2^2-4(2)+3 so f(2) = 4-8 +3 so f(2) = -1. alright, so that's the y coordinate?

OpenStudy (anonymous):

yes, the y-coordinate of the vertex the vertex is the point (2,-1)

OpenStudy (anonymous):

ok. so let me just review here. the x coordinate to the vertex is -b/2a, which we then plug in for f(x) and solve for the y coordinate. where do the two side points come from?

OpenStudy (anonymous):

what 2 side points?

OpenStudy (anonymous):

the vertex is the middle one, but there have to be two other points, don't there? it's not a parabola with one point

OpenStudy (anonymous):

there are an infinite number of points on a parabola, not just 3

OpenStudy (anonymous):

you want the x-intercepts, the y-intercept, the vertex, and the axis of symmetry. a few other points won't hurt either

OpenStudy (anonymous):

yeah, but you are only require to graph 3... so where do i get the other 2? :o

OpenStudy (anonymous):

you already found the x-intercepts, (3, 0) and (1, 0)

OpenStudy (anonymous):

the y-intercept is (0, 3) not y=3

OpenStudy (anonymous):

ohhhh ok. so the three points are (3,0), (1,0), and (2, -1)

OpenStudy (anonymous):

you could use those, i'd also use a 4th point for the y-intercept of (0,3)

OpenStudy (anonymous):

oh ok. thank you sooooo much. you're a life saver!

OpenStudy (anonymous):

yw